# Entanglement, Spin Squeezing and Quantum Sensing in a Spin-5/2 Heisenberg Molecular Iron(III) Triangle

### v2

@@ V.1 Sequential measurement
The sequential measurement protocol originally demonstrated in Ref. Montenegro2022 offers an innovative approach to optimize sensing efficiency by exploiting measurement-induced dynamics. According to the protocol’s description given in Ref. Montenegro2022 , let us outline the key steps of this protocol in the context of single-parameter estimation when specifically focusing on estimating a local trasnverse magnetic field $B_{x}$ in absence of the longitudinal magnetic field $B_{z}=0$. First, the system is initialized in one of four lowest Dicke states $|D_{0}\rangle$, $|D_{1}\rangle$, $|D_{2}\rangle$, or $|D_{3}\rangle$. Then, the system undergoes the following sequential measurements:
@@
This protocol efficiently utilizes time resources by avoiding the overhead time required for reinitialization after each measurement. Instead, a sequence of measurements is performed successively on the same probe with intermittent-free evolution periods. This strategy allows for effective data gathering while minimizing initialization overhead. Further details of this protocol can be found in the original work Montenegro2022 and its subsequent studies Montenegro2023 ; Montenegro2024 .
V.2 Parameter estimation
Let us now proceed with calculating the classical Fisher information (CFI) $F(B_{x})$ Yuan2020 , which will be used for estimating a local magnetic field $B_{x}$ using the measurement basis of the third spin as described by the conventional precision bound. The CFI is a concept from classical statistics that quantifies the amount of information that an observable random variable provides about an unknown parameter. It measures the sensitivity of the likelihood function to changes in the parameter. In our case, the CFI is given by Montenegro2022 :
@@ V.2 Parameter estimation
Figure 7: The inverse of classical Fisher information $F^{-1}$ versus the number of sequential measurements $n_{\mathrm{seq}}$ performed at the third spin as calculated according to Eq. (11) for the spin-5/2 Heisenberg triangle with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$. This parameter set was adjusted for a theoretical modeling of the molecular complex Fe3 Plass2023 initialized in one of four Dicke states: (a), (b) $|D_{0}\rangle$; (c), (d) $|D_{1}\rangle$; (e), (f) $|D_{2}\rangle$; (g), (h) $|D_{3}\rangle$. The left and right panels show the respective results for $F^{-1}$ at two selected values of the magnetic field $B_{x}=1\mathrm{T}$ and $B_{x}=5\mathrm{T}$, respectively. The fitting functions $g(n_{\mathrm{seq}})=\alpha n_{\mathrm{seq}}^{-\beta}$ with the exponent coefficient $\beta>1$ are also shown.
@@
To observe the effect of sequential measurements on sensing precision of the molecular compound Fe3, we present in Fig. 7 the inverse of the CFI $F^{-1}$ as a function of the number of sequential measurements $n_{\mathrm{seq}}$ performed at the third spin. The panels compare results for different initial states and two selected magnetic-field strengths. For example, Fig. 7(a) and (b) shows $F^{-1}$ for the initial ferromagnetic Dicke state $|D_{0}\rangle$ by considering the local magnetic field $B_{x}=1\mathrm{T}$. and $B_{x}=5\mathrm{T}$, respectively. Evidently, the inverse of CFI in Figs. 7(a) and 7(b) decreases rapidly with increasing the number of sequential measurements $n_{\mathrm{seq}}$. This observation would suggest that the precision of sensing improves significantly when performing greater number of sequential measurements. The fitting function $g(n_{\mathrm{seq}})=\alpha n_{\mathrm{seq}}^{-\beta}+\epsilon$ with $\epsilon$ being negligible and $\beta$ consistently greater than 1, suggests that the sequential measurements enable a quantum-enhanced sensitivity beyond the classical limit Montenegro2022 .
@@
Figs. 7(c)-7(h) display the inverse of CFI for more complex initial Dicke states $|D_{1}\rangle$, $|D_{2}\rangle$, or $|D_{3}\rangle$ with character of a quantum superposition of states by considering two different local magnetic fields $B_{x}=1\mathrm{T}$ (left panel) and $B_{x}=5\mathrm{T}$ (right panel). It is evident from Figs. 7(c)-7(h) that the inverse of CFI decreases significantly as the number of sequential measurements $n_{\mathrm{seq}}$ increases. This behavior supports the conclusion that sequential measurements using the trinuclear molecular complex Fe3 substantially enhance sensing precision regardless of the complexity of the initial Dicke state with the nature of quantum superposition of states. The decay of inverse CFI with increasing the number of sequential measurements can be satisfactorily fitted with a power-law dependence with the exponent $\beta>1$, which in turn confirms the advantage of sequential measurements for the high-precision sensing of the magnetic field.
VI Conclusions
@@
Our findings demonstrate that the molecular complex Fe3 exhibits sizable bipartite and tripartite entanglement, which are thermally quite robust and persist up to temperatures around 30 K and 70 K, respectively. We have additionally examined coherent, squeezed, and unsqueezed states within the molecular compound Fe3. The squeezed states, which are particularly relevant for applications in quantum technologies, can be realized at temperatures below 15 K and magnetic fields under 25 T. When a local magnetic field is applied to the first spin of the molecular complex Fe3 and the local magnetization of the third spin is measured, a remarkable synchronization is detected between the readout and sensor spin. This synchronization between the spin dynamics of the readout and sensor spin suggests that the measurement of the third spin can provide remote information about the local field at the first spin, which makes the molecular complex Fe3 an effective magnetic sensor for quantum information processing.
@@
The outcomes of this work could inspire further research on other molecular magnets, more specifically, in the area of quantized magnetization plateaus, entanglement, spin squeezing, and quantum sensing in various transition-metal complexes. This comprehensive exploration provides valuable insights into several intriguing quantum phenomena exhibited by the molecular complex Fe3 and highlights the potential of related molecular quantum magnets for near-term applications in quantum technologies.
@@ Acknowledgments
$|3/27\rangle^{\mathrm{ii}}$ $=$ $0.26|5/2,3/2,-5/2\rangle-0.20|5/2,1/2,-3/2\rangle+0.15|5/2,-1/2,-1/2\rangle$

### v3

@@ IV.1 Sequential measurement
The sequential measurement protocol originally demonstrated in Ref. [54] offers an innovative approach to optimize sensing efficiency by exploiting measurement-induced dynamics. According to the protocol’s description given in Ref. [54], let us outline the key steps of this protocol in the context of single-parameter estimation when specifically focusing on estimating a local trasnverse magnetic field $B_{x}$ in absence of the longitudinal magnetic field $B_{z}=0$. First, the system is initialized in one of four lowest Dicke states $|D_{0}\rangle$, $|D_{1}\rangle$, $|D_{2}\rangle$, or $|D_{3}\rangle$. Then, the system undergoes the following sequential measurements:
@@
This protocol efficiently utilizes time resources by avoiding the overhead time required for reinitialization after each measurement. Instead, a sequence of measurements is performed successively on the same probe with intermittent-free evolution periods. This strategy allows for effective data gathering while minimizing initialization overhead. Further details of this protocol can be found in the original work [54] and its subsequent studies [55, 56].
IV.2 Parameter estimation
Let us now proceed with calculating the classical Fisher information (CFI) $F(B_{x})$ [81], which will be used for estimating a local magnetic field $B_{x}$ using the measurement basis of the third spin as described by the conventional precision bound. The CFI is a concept from classical statistics that quantifies the amount of information that an observable random variable provides about an unknown parameter. It measures the sensitivity of the likelihood function to changes in the parameter. In our case, the CFI is given by [54]:
@@ IV.2 Parameter estimation
Figure 6: The inverse of classical Fisher information $F^{-1}$ versus the number of sequential measurements $n_{\mathrm{seq}}$ performed at the third spin as calculated according to Eq. (8) for the spin-5/2 Heisenberg triangle with the coupling constant $J=12.56\,\mathrm{cm}^{-1}$ and the gyromagnetic factor $g=2.0$. This parameter set was adjusted for a theoretical modeling of the molecular complex Fe3 [72] initialized in one of four Dicke states: (a), (b) $|D_{0}\rangle$; (c), (d) $|D_{1}\rangle$; (e), (f) $|D_{2}\rangle$; (g), (h) $|D_{3}\rangle$. The left and right panels show the respective results for $F^{-1}$ at two selected values of the magnetic field $B_{x}=1\mathrm{T}$ and $B_{x}=5\mathrm{T}$, respectively. The fitting functions $g(n_{\mathrm{seq}})=\alpha n_{\mathrm{seq}}^{-\beta}$ with the exponent coefficient $\beta>1$ are also shown.
@@
To observe the effect of sequential measurements on sensing precision of the molecular compound Fe3, we present in Fig. 6 the inverse of the CFI $F^{-1}$ as a function of the number of sequential measurements $n_{\mathrm{seq}}$ performed at the third spin. The panels compare results for different initial states and two selected magnetic-field strengths. For example, Fig. 6(a) and (b) shows $F^{-1}$ for the initial ferromagnetic Dicke state $|D_{0}\rangle$ by considering the local magnetic field $B_{x}=1\mathrm{T}$. and $B_{x}=5\mathrm{T}$, respectively. Evidently, the inverse of CFI in Figs. 6(a) and 6(b) decreases rapidly with increasing the number of sequential measurements $n_{\mathrm{seq}}$. This observation would suggest that the precision of sensing improves significantly when performing greater number of sequential measurements. The fitting function $g(n_{\mathrm{seq}})=\alpha n_{\mathrm{seq}}^{-\beta}+\epsilon$ with $\epsilon$ being negligible and $\beta$ consistently greater than 1, suggests that the sequential measurements enable a quantum-enhanced sensitivity beyond the classical limit [54].
@@
Figs. 6(c)-6(h) display the inverse of CFI for more complex initial Dicke states $|D_{1}\rangle$, $|D_{2}\rangle$, or $|D_{3}\rangle$ with character of a quantum superposition of states by considering two different local magnetic fields $B_{x}=1\mathrm{T}$ (left panel) and $B_{x}=5\mathrm{T}$ (right panel). It is evident from Figs. 6(c)-6(h) that the inverse of CFI decreases significantly as the number of sequential measurements $n_{\mathrm{seq}}$ increases. This behavior supports the conclusion that sequential measurements using the trinuclear molecular complex Fe3 substantially enhance sensing precision regardless of the complexity of the initial Dicke state with the nature of quantum superposition of states. The decay of inverse CFI with increasing the number of sequential measurements can be satisfactorily fitted with a power-law dependence with the exponent $\beta>1$, which in turn confirms the advantage of sequential measurements for the high-precision sensing of the magnetic field.
V Conclusions
@@
Our findings demonstrate that the molecular complex Fe3 exhibits sizable bipartite and tripartite entanglement, which are thermally quite robust and persist up to temperatures around 30 K and 70 K, respectively. In addition, we have found that a remarkable synchronization is detected between the readout and sensor spin when a local magnetic field is applied to the first spin of the molecular complex Fe3 and the local magnetization of the third spin is measured. This synchronization between the spin dynamics of the readout and sensor spin suggests that the measurement of the third spin can provide remote information about the local field at the first spin, which makes the molecular complex Fe3 an effective magnetic sensor for quantum information processing.
@@
The outcomes of this work could inspire further research on other molecular magnets, more specifically, in the area of quantized magnetization plateaus, entanglement, and quantum sensing in various transition-metal complexes. This comprehensive exploration provides valuable insights into several intriguing quantum phenomena exhibited by the molecular complex Fe3 and highlights the potential of related molecular quantum magnets for near-term applications in quantum technologies.
@@ Acknowledgments
$|3/2\rangle^{\mathrm{ii}}$ $=$ $0.26|5/2,3/2,-5/2\rangle-0.20|5/2,1/2,-3/2\rangle+0.15|5/2,-1/2,-1/2\rangle$